OpenStudy (anonymous):

4sqrt(1-((x^2)/16))=sqrt(16-x^2) why?

OpenStudy (jusaquikie):

4=\[4=\sqrt{16}\]

geerky42 (geerky42):

\[4 = \sqrt{16}\] So \(4\sqrt{1-\dfrac{x^2}{16}} = \sqrt{16 \left( 1-\dfrac{x^2}{16} \right)} = \boxed{ \sqrt{16 - x^2}}\)